[Brahms-l] Shift report 20050614 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 14 2005 - 23:45:33 EDT
Shifter leader was BB
Shifters were AG


1. Air conditioning in FEH fixed ~5:30 pm<br>
   Electronics in FEH reset ~6 pm, with<br>
   help from Bob and Hiro<br>

2. Data runs 15483-15487 (~600k trigger 5)<br>

3. Pedestal runs<br>

4. D1 and D2 tripped several times<br>

5. Beam was available for ~6 hours<br>

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 14 23:45:50 2005

This archive was generated by hypermail 2.1.8 : Tue Jun 14 2005 - 23:46:02 EDT